solution set for
2(x-1)less than or equal to 4x-2

Answer by checkley77(12844)

2(x-1)<=4x-2
2x-2<=4x-2
2x-4x<=-2+2
-2x<=0
x>=0/-2 when dividing by a negative number the < sign changes to a > sign.
x>=0
Proof:
Let x=1
2(1-1)<=4*1-2
2*0<=4-2
0<=2